# Towards a Sharp Analysis of Offline Policy Learning for $f$-Divergence-Regularized Contextual Bandits

### v2

@@ 3.2 Algorithm and Main Results
which enables us to bound $({g^{*}}-{\bar{g}})^{\top}H^{*}(\widetilde{g})({g^{*}}-{\bar{g}})$ by $\alpha^{-1}\eta\mathbb{E}_{\pi^{\mathsf{ref}}}[({g^{*}}-{\widehat{g}})^{2}]$. Since $\mathbb{E}_{\pi^{\mathsf{ref}}}[({g^{*}}-{\widehat{g}})^{2}]$ is not related to the optimal $\pi^{*}$, our resulting upper bound does not depend on any notion of concentrability.
4 Generalization to Contextual Dueling Bandits
In this section, we extend our algorithm to the problems of regularized contextual dueling bandits, where the learner receives preference comparison instead of absolute signals. Our data model follows Zhu et al. (2023); Zhan et al. (2023) and the notion of suboptimality follows Xiong et al. (2024); Zhao et al. (2024).
4.1 Problem Setup
We still consider contextual bandits $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$ where ${\mathcal{S}}$ is the state space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function. [^1] We overload some notations in Section 2 by their dueling counterparts for notational simplicity. But only relative preference feedback is available, viz., we have an $\mathrm{i.i.d.}$ offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$, where $s_{i}\in{\mathcal{S}}$ is generated from distribution $\rho$ and $a^{1}_{i},a^{2}_{i}\sim\pi^{\mathsf{ref}}$. The binary preference label $y_{i}=1$ indicates $a_{i}^{1}$ is preferred over $a_{i}^{2}$ (denoted by $a^{1}\succ a^{2}$) and 00 for $a^{2}\succ a^{1}$ given context $s$. In this work we consider the Bradley-Terry Model, where $\mathbb{P}[y=1|s,a^{1},a^{2}]=\sigma(r(s_{i},a^{1}_{i})-r(s_{i},a^{2}_{i}))$, where $\sigma(x)=(1+e^{-x})^{-1}$ is the link function. The objective here identical to (2.1) for KL-regularization and (3.1) for $f$-divergence regularization. Our goal is still to find an $\epsilon$-optimal policy. To control the complexity of the function class ${\mathcal{G}}$, we assume that Assumption 2.1 still holds here.
Concentrability.
Definition 4.1.
Assumption 4.2 (All-policy concentrability).
Assumption 4.3 (Single-policy concentrability).
Similar single-policy concentrability assumptions have appeared in previous work in offline contextual dueling bandits (Huang et al., 2024; Song et al., 2024) and similar notions has also appeared in the analysis of model-based RL (Uehara and Sun, 2021; Wang et al., 2024). Although Assumption 4.3 is strictly weaker than Assumption 4.2, $D^{2}_{\pi^{*}}\leq|{\mathcal{S}}||\mathcal{A}|C^{\pi^{*}}$ still holds and, in general, $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ cannot be bounded by each other up to some constant factors.
4.2 Main Results
In this subsection, we provide our main results for contextual dueling bandits with regularization. We refer the reader to Appendix D for a more detailed discussion. We first present the following results for KL-regularized contextual dueling bandit.
Theorem 4.4 (Informal).
There exists an algorithm, with high probability, generates $\epsilon$-optimal policy with $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon))$ samples. Moreover, for any sufficiently small $\epsilon>0$, and $\eta>0$, and any algorithm Alg, there is a KL-regularized contextual dueling bandit such that Alg requires at least $\Omega\big{(}\min\{{\eta C^{\pi^{*}}\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)},{(C^{\pi^{*}})^{2}\epsilon^{-2}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
The following theorem shows a similar guarantee for $f$-divergence regularized contextual dueling bandits where $f$ is $\alpha$-strongly convex.
Theorem 4.5 (Informal).
Under Assumption 3.1, there exists an algorithm generating $\epsilon$-optimal policy with $n=\widetilde{O}(\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon))$ with high probability. Moreover, for any sufficiently small $\epsilon>0$, $\eta,\alpha>0$, and any offline RL algorithm Alg, there is an $\alpha$-strongly convex function $f$ such that there exists an $f$-divergence regularized contextual dueling bandit such that Alg requires at least $\Omega\big{(}{\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)}\big{)}$ samples to return an $\epsilon$-optimal policy.
Theorems 4.4 and 4.5 are extension of Theorems 2.9 and 3.2 to offline contextual dueling bandits. These theorems indicate that, for both KL-regularization and $f$-divergence regularization, our techniques can be adapted to achieve algorithms and achieve similar sample complexities.
5 Conclusion and Limitation
@@
In this work, we take the first step towards fully understanding the statistical efficiency with respect to $f$-divergence-regularized objectives of offline policy learning by sharp analyses for two empirically relevant subclasses. (1) We are the first to show that single-policy concentrability is nearly the right coverage condition for reverse KL to achieve the fast $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity. The novel techniques in algorithm analysis leverages the curvature of KL-regularized objectives and integrates pessimism with a newly identified moment-based observation, enabling a neat refinement of a mean-value-type argument to the extreme; which are decoupled from tricky algorithmic tweaks, and thus might be of independent interest. (2) If strong convexity is further imposed on $f$, our fast $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity is provably free of any coverage dependency. Unlike those for KL, the upper bound arguments for strongly convex $f$ do not rely on the specific closed-form of the regularized objective minimizer.
@@
All techniques in this work can be generalized beyond vanilla absolute reward feedback, as certified by contextual dueling bandits. However, there is still a gap between the upper bound and lower bound under reverse-KL regularization. Also, for general $f$-divergence other than reverse-KL, our results are limited to twice continuously differentiable and strongly convex $f$. The fully closing of the gap under reverse-KL regularization and the investigation of general $f$-divergence are left as future directions.
Appendix A Review of Existing Results
@@
Besides the notation introduced in Section 1, we will use the following notations in Appendix. We denote $[N]\coloneqq\{1,\cdots,N\}$ for any positive integer $N$. Boldfaced lower case (resp. upper case) letters are reserved for vectors (resp. matrices). Given a positive definite $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ and $\mathbf{x}\in\mathbb{R}^{d}$, we denote the vector’s Euclidean norm by $\|\mathbf{x}\|_{2}$ and define $\|\mathbf{x}\|_{\bm{\Sigma}}=\sqrt{\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}}$. We use $\mathsf{Bern}(p)$ to denote Bernoulli distribution with expectation $p$ and $\mathsf{Unif}(\mathcal{X})$ for the uniform distribution on finite set $\mathcal{X}$. For $x\in\mathbb{R}^{\mathcal{A}}$, we denote $\|x\|_{1}=\sum_{a\in\mathcal{A}}|x_{a}|$. We also denote $x_{n}=\Omega(y_{n})$ by $x_{n}\gtrsim y_{n}$ in Appendix. We use $d_{H}$ for Hamming distance.
@@ A.1 Previous Attempts on Understanding KL-regularized RL
Several recent studies (Xie et al., 2024; Xiong et al., 2024; Zhao et al., 2024, 2025; Foster et al., 2025) switched the focus to analyzing the suboptimality guarantee with respect to the regularized objective (1.1). In particular, Xie et al. (2024) studied token-level Markov decision processes (MDPs) and proposed a KL-regularized RL algorithm named XPO, which achieves $\widetilde{O}(\epsilon^{-2})$ sample complexity under their notion of all-policy concentrability. Xiong et al. (2024) proposed an Offline GSHF algorithm via the principle of pessimism in the face of uncertainty, and proved $\widetilde{O}(\epsilon^{-2})$ sample complexity under single-policy concentrability (See Section 2.1 for detailed definitions of concentrability). On the other hand, the sharp analysis in Zhao et al. (2024) yields the optimal sample complexity $\widetilde{O}(\epsilon^{-1})$, but requires all-policy concentrability (Zhao et al., 2024, Definition 2.6), i.e., the behavior policy $\pi^{\mathsf{ref}}$ is required to cover the entire function class for all possible policies. Zhao et al. (2025) considered the online episodic MDP setting, which inherently does not need any notion of data coverage and thus their results are not directly adaptable to our offline setting. Foster et al. (2025) considered an interesting hybrid setting in which the $n$ state-action pairs are still from the offline dataset but $\Omega(n)$ online reward queries and policy switches are allowed; in contrast, in our setting, all reward signals are obtained in a purely offline fashion.
Here, we briefly discuss the direct adaptation of previous sample complexity results (with respect to KL-regularized objectives) to our setting and demonstrate the reason why theirs cannot imply an $\widetilde{O}(\epsilon^{-1})$ sample complexity without all-policy concentrability. Specifically, Xiong et al. (2024) obtained a performance gap upper bound under linear function approximation
@@
where $\lambda_{\min}$ and $\lambda_{\max}$ is the minimum and maximum eigenvalue of a matrix, the first inequality holds due to the fact that $\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}\geq\|\mathbf{x}\|_{2}^{2}\lambda_{\min}(\bm{\Sigma})$ and the second inequality holds due to $\lambda_{\max}(\bm{\Sigma})\leq\text{tr}(\bm{\Sigma})$. Zhao et al. (2024) proposed a two-stage learning algorithm and obtained an $\widetilde{O}(\epsilon^{-1})$ sample complexity for online KL-regularized bandits. The algorithm can be adopted to offline learning by removing the second stage [^2] This can be done by setting the $n$ in their paper to 00. and treat the samples from first stage as the offline dataset. An analogous analysis gives a sample complexity of $\widetilde{O}(D^{2}\epsilon^{-1})$, where $D^{2}$ is the all-policy concentrability.
Appendix B Missing Proofs from Section 2
B.1 Proof of Lemma 2.12

### v3

@@ 3.2 Algorithm and Main Results
which enables us to bound $({g^{*}}-{\bar{g}})^{\top}\nabla^{2}H^{*}(\widetilde{g})({g^{*}}-{\bar{g}})$ by $\alpha^{-1}\eta\mathbb{E}_{\pi^{\mathsf{ref}}}[({g^{*}}-{\widehat{g}})^{2}]$. Since $\mathbb{E}_{\pi^{\mathsf{ref}}}[({g^{*}}-{\widehat{g}})^{2}]$ is not related to $\pi^{*}$, the upper bound is independent of any notion of concentrability.
(a) The empirical relation between $\log_{2}n$ and $\log_{2}\mathrm{SubOpt}$. The fitted rate means the slope of $\log_{2}n\sim\log_{2}\mathrm{SubOpt}$ estimated via linear regression. Here $n$ is the sample size. Every point is the average over 100 independent trials.
4 Experiments
Simulation on multi-armed bandits. We first empirically check the correctness of our matching bounds for KL and $f$-divergence on the simplest testbed: two-armed bandits, i.e., $\mathcal{A}=\{0,1\}$. We use one hard instance constructed in the proof of Theorem 2.11 (Section D.5) for the simulation under KL and one hard instance constructed in the proof of Theorem 3.4 (Section E.2) for the simulation under $f$-divergence with $f(x)=\alpha(x-1)^{2}/2$. Recall that the dependency on $\epsilon$ in all sample complexity bounds above is $\widetilde{\Theta}(\epsilon^{-1})$, and thus both $\mathrm{SubOpt}_{\mathrm{RKL}}$ and $\mathrm{SubOpt}_{f\mathrm{div}}$ should be roughly proportional to $n^{-1}$ as a function of the sample size $n$, which can be verified from the linear regression between $\log_{2}n$ and $\log_{2}\mathrm{SubOpt}$; i.e., the estimated slope should be approximately 1-1. Therefore, the two fitted rates in LABEL:subfig:kl indicates that KL-PCB indeed achieves the near-optimal statistical rate $n^{-1}$ under different $\pi^{\mathsf{ref}}$’s and the counterparts in the LHS of LABEL:subfig:f indicates the near-optimality of $f$-CB empirically. The contrast between LABEL:subfig:kl and the LHS of LABEL:subfig:f also corroborates that the sample complexity against the KL-regularized objective positively depend on the concentrability, while that against the $\chi^{2}$-divergence-regularized objective does not vary with the coverage condition of $\pi^{\mathsf{ref}}$. Moreover, on top of the hard instance for $f$-divergence, we further set $\alpha=2^{15}/n$ to numerically examine the scaling of $\mathrm{SubOpt}_{f\mathrm{div}}$ w.r.t. the strong convexity modulus $\alpha$. As shown on the RHS of LABEL:subfig:f, $\mathrm{SubOpt}_{f\mathrm{div}}$ remains stable as $n$ goes up given $n\alpha\equiv 2^{15}$; therefore, LABEL:subfig:f also empirically verified that $\mathrm{SubOpt}_{f\mathrm{div}}$ is inversely proportional to $\alpha$.
(b) The empirical relation between $\log_{2}n$ and $\log_{2}\mathrm{SubOpt}$ for linear bandits. In the legend, we denote $C^{\pi^{*}}$ (resp. $D^{2}_{\pi^{*}}$) by C (resp. Dˆ2).
Simulation on linear bandits. We then simulate a linear bandit as follows. The constructions of the feature map $\bm{\phi}$, ground-truth parameter $\bm{\theta}^{*}$ and the induced reward are detailed in Section C.1. The behavior policy is constructed as $\pi^{\mathsf{ref}}=\beta\mathsf{Unif}(\mathcal{A})+(1-\beta)\mathsf{Unif}(\mathcal{A}_{k})$, where $\mathcal{A}_{k}\subset\mathcal{A}$ be the subset such that $\mathcal{A}_{k}$ consists of the $k$ arms with the lowest expected reward. We consider three different behavior policies, $(\beta,k)\in\{(1,\cdot),(0.1,4),(0.05,20)\}$, which induces various $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ so as to demonstrate the influence of coverage under different regularization. The results are compiled in Figure 2(b). Specifically, for the KL-regularized cases depicted in Figure LABEL:subfig:kl-linear, we see that as the coverage coefficients $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ vary, there is a consistent sub-optimality gap margin between these instances. In contrast, LABEL:subfig:chi2-linearshows that the sub-optimality gaps under different instances (with distinct coverage coefficient) are very close for sufficiently large sample sizes. These results corroborate our theoretical finding that the sample complexity w.r.t. KL-regularized objectives is concentrability-dependent but that w.r.t $f$-divergence ones is not (for strongly convex $f$).
(c) The empirical relation between $\log_{2}n$ and $\log_{2}\mathrm{SubOpt}$ on MNIST dataset.
Real-world experiments. We further verify our theory on a vision dataset, MNIST (LeCun, 1998). The construction of the feature map is detailed in Section C.2. We consider two reference polices, a uniform policy $\mathsf{Unif}(\mathcal{A})$ and a spiky policy $0.5\mathsf{Unif}(\mathcal{A})+0.5\mathsf{Dirac}(\{0\})$ to obtain instances with different concentrability coefficients. Figure 2(c) exhibits the $\mathrm{SubOpt}$ curves, which show that under KL-regularization, when sample size is not large enough, there exists a considerable gap between instances with different behavior policy, but the gap is vanishing as the sample size increases. On the other hand, as for $\chi^{2}$-divergence regularization, such a gap vanishes quickly when the sample size becomes moderately large and the sub-optimal gap remains similar for larger sample sizes. These results are consistent with the simulation in Section 4 and our theoretical findings.
5 Conclusion and Future Work
@@
In this work, we take the first step towards fully understanding the statistical efficiency with respect to $f$-divergence-regularized objectives of offline policy learning by sharp analyses for two empirically relevant subclasses. (1) We are the first to show that single-policy concentrability is nearly the right coverage condition for reverse KL to achieve the fast $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity. The novel techniques in algorithm analysis leverages the curvature of KL-regularized objectives and integrates pessimism with a newly identified moment-based observation, enabling a neat refinement of a mean-value-type argument to the extreme; which are decoupled from tricky algorithmic tweaks, and thus might be of independent interest. (2) If strong convexity is further imposed on $f$, our fast $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity is provably free of any coverage dependency. Unlike those for KL, the upper bound arguments for strongly convex $f$ do not rely on specific closed-form solutions of the regularized objective maximizer.
@@
All techniques in this work can be generalized beyond vanilla absolute reward feedback, as certified by CDBs, which is detailed in Appendix F under a slightly different notion of $D^{2}$ tailored for pairwise comparison feedback. However, for reverse-KL regularization, the $D^{2}_{\pi^{*}}$ in the upper bound and the $C^{\pi^{*}}$ in the lower bound still does not perfectly match. Also, for general $f$-divergence other than reverse-KL, our analyses require $f$ to be twice-continuously differentiable and strongly convex. Fully closing the gap under reverse-KL regularization and extending the analysis to general $f$-divergences are interesting directions for future work.
The Use of Large Language Models (LLMs)
We use LLMs as a tool to refine our writing and correct grammatical errors.
Acknowledgment
We thank the anonymous reviewers and area chair for their helpful comments. QZ, KJ, HZ and QG are supported in part by the National Science Foundation DMS-2323113 and IIS-2403400. HZ is also supported in part by Amazon PhD Fellowship. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
Appendix A Additional Review of Existing Results
@@
Besides the notation introduced in Section 1, we will use the following notations in Appendix. We denote $N]\coloneqq\{1,\cdots,N\}$ for any positive integer $N$. Boldfaced lower case (resp. upper case) letters are reserved for vectors (resp. matrices). Given a positive definite $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ and $\mathbf{x}\in\mathbb{R}^{d}$, we denote the vector’s Euclidean norm by $\|\mathbf{x}\|_{2}$ and define $\|\mathbf{x}\|_{\bm{\Sigma}}=\sqrt{\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}}$. We use $\mathsf{Bern}(p)$ to denote Bernoulli distribution with expectation $p$ and $\mathsf{Unif}(\mathcal{X})$ for the uniform distribution on finite set $\mathcal{X}$. For $x\in\mathbb{R}^{|\mathcal{A}|}$, we denote $\|x\|_{1}=\sum_{a\in\mathcal{A}}|x_{a}|$. We also denote $x_{n}=\Omega(y_{n})$ by $x_{n}\gtrsim y_{n}$ in Appendix. We use $d_{H}$ for Hamming distance.
@@ A.1 Previous Attempts on Understanding KL-regularized RL
Several recent studies (Xie et al., 2024; Xiong et al., 2024; Zhao et al., 2024; 2025; Foster et al., 2025) switched the focus to analyzing the sub-optimality guarantee with respect to the regularized objective (1.1). In particular, Xie et al. (2024) studied token-level Markov decision processes (MDPs) and proposed a KL-regularized RL algorithm named XPO, which achieves $\widetilde{O}(\epsilon^{-2})$ sample complexity under their notion of all-policy concentrability. Xiong et al. (2024) proposed an Offline GSHF algorithm via the principle of pessimism in the face of uncertainty, and proved $\widetilde{O}(\epsilon^{-2})$ sample complexity under single-policy concentrability (See Section 2.1 for detailed definitions of concentrability). On the other hand, the sharp analysis in Zhao et al. (2024) yields the optimal sample complexity $\widetilde{O}(\epsilon^{-1})$, but requires all-policy concentrability (Zhao et al., 2024, Definition 2.6), i.e., the behavior policy $\pi^{\mathsf{ref}}$ is required to cover the entire function class for all possible policies. Zhao et al. (2025) considered the online episodic MDP setting, which inherently does not need any notion of data coverage and thus their results are not directly adaptable to our offline setting. Foster et al. (2025) considered an interesting hybrid setting in which the $n$ state-action pairs are still from the offline dataset but $\Omega(n)$ online reward queries and policy switches are allowed; in contrast, in our setting, all reward signals are obtained in a purely offline manner.
Previous analyses and results in detail.
Here, we briefly discuss the direct adaptation of previous sample complexity analysis and results (with respect to KL-regularized objectives) to our setting and demonstrate the reason why theirs cannot imply an $\widetilde{O}(\epsilon^{-1})$ sample complexity without all-policy concentrability. In previous analysis of pessimism for unregularized objectives (Jin et al., 2021; Xiong et al., 2022), the sub-optimality gap is decomposed via the performance difference lemma as follows
[…]
$\leq\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{\widehat{g}}(a)]-\eta^{-1}\mathsf{KL}(\pi^{*}\|\pi^{\mathsf{ref}})+\eta^{-1}\mathsf{KL}(\widehat{\pi}\|\pi^{\mathsf{ref}})$
$\leq\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(a)]-\mathbb{E}_{a\sim\pi^{*}}[{\widehat{g}}(a)]-\eta^{-1}\mathsf{KL}(\pi^{*}\|\pi^{\mathsf{ref}})+\eta^{-1}\mathsf{KL}(\pi^{*}\|\pi^{\mathsf{ref}})$
$=\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(a)-{\widehat{g}}(a)],$
where the first inequality holds due to pessimism and last inequality holds due to $\widehat{\pi}$ is optimal for $^{\widehat{g}}$. Notably, the KL-regularization term is canceled out in the analysis, leading to a loose sample complexity $\widetilde{O}(\epsilon^{-2})$ since the curvature of KL-divergence is not exploited. Specifically, under linear function approximation, this performance gap, obtained by Xiong et al. (2024) becomes
@@
where $\lambda_{\min}$ and $\lambda_{\max}$ is the minimum and maximum eigenvalue of a matrix, the first inequality holds due to the fact that $\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}\geq\|\mathbf{x}\|_{2}^{2}\lambda_{\min}(\bm{\Sigma})$ and the second inequality holds due to $\lambda_{\max}(\bm{\Sigma})\leq\text{tr}(\bm{\Sigma})$. Zhao et al. (2024) proposed a two-stage learning algorithm and obtained an $\widetilde{O}(\epsilon^{-1})$ sample complexity for online KL-regularized bandits. The algorithm can be adopted to offline learning by removing the second stage [^5] This can be done by setting the $n$ in their paper to 0. and treat the samples from first stage as the offline dataset. An analogous analysis gives a sample complexity of $\widetilde{O}(D^{2}\epsilon^{-1})$, where $D^{2}$ is the all-policy concentrability.
Appendix B Additional Discussion of Relation between Coverage Measures
In this section, we provide more illustrations on the relation between two coverage measures, $D^{2}_{\pi^{*}}$ and $C^{\pi^{*}}$. In particular, we provide two cases under linear function approximation, on one of which $D^{2}_{\pi^{*}}=\Theta(dC^{\pi^{*}})$ and on the other we have $D^{2}_{\pi^{*}}\ll C^{\pi^{*}}$, where $d$ is the dimension of the function class. We summarized them as two propositions.
Proposition B.1.
There exist a KL-regularized linear bandit instance, such that $D^{2}_{\pi^{*}}=\Theta(dC^{\pi^{*}})$.
Proof.
We construct the instance as follows. Let $d=2A+1$ be some odd number and consider an $2A+1$-armed bandit, such that the feature vector of the $i$-th arm, $\bm{\phi}(a_{i})=\mathbf{e}_{i}\in\mathbb{R}^{d}$, which has $1$ on its $i$-th entry and 0 on all other entries. The reference policy $\pi^{\mathsf{ref}}(a_{i})=(2AC)^{-1}$ for $i\in[2A]$ and $\pi^{\mathsf{ref}}(a_{2A+1})=(C-1)/C$, where $2C-1=e^{\eta}$. The ground truth reward function $\bm{\theta}^{*}=\sum_{i\leq A}\mathbf{e}_{i}$ and the function class is given by all $\|\bm{\theta}\|_{\infty}\leq 1$. By construction, we know that $\pi^{*}(a_{i})\geq\pi^{\mathsf{ref}}(a_{i})$ if and only if $i\in[A]$ and its closed form is given by
$\pi^{*}(a_{i})=\frac{1}{A}\frac{e^{\eta}}{e^{\eta}+2C-1}=\frac{1}{2A},$
which gives $C^{\pi^{*}}=C$. Now we compute the $D^{2}_{\pi^{*}}$ of this instance. For all $i\in[A]$, we know that
$D^{2}(a_{i})=\sup_{\|\bm{\theta}\|_{\infty}\leq 2}\frac{\left\langle\bm{\theta},\mathbf{e}_{i}\right\rangle^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\left\langle\bm{\theta},\mathbf{e}_{j}\right\rangle^{2}}=2CA=\Theta(Cd),$
where the second equation holds with $\bm{\theta}=\mathbf{e}_{i}$. Taking expectation over $\pi^{*}$, we have
$D^{2}_{\pi^{*}}\geq\sum_{i\in[A]}D^{2}(a_{i})=\Theta(C^{\pi^{*}}d),$
which concludes the proof. ∎
The following proposition provides another instance on which $D^{2}_{\pi^{*}}\ll C^{\pi^{*}}$.
Proposition B.2.
For any $C\geq 2$, there exists a KL-regularized linear bandit instance, such that $C^{\pi^{*}}=C/2$ and $D^{2}_{\pi^{*}}=\Theta(1)$.
Proof.
We consider the function class of $\bm{\theta}\in\mathbb{R}^{2}$ and $\|\bm{\theta}\|\leq\sqrt{2}$. The instance consists of three arms, where $\bm{\phi}(a_{1})=(1,0)$, $\bm{\phi}(a_{2})=(0,1)$, and $\bm{\phi}(a_{3})=(1,1)$. The ground truth parameter $\bm{\theta}^{*}=(1,1)$. The reference policy is given by $\pi^{\mathsf{ref}}(a_{1})=\pi^{\mathsf{ref}}(a_{2})=1/2-1/2C$ and $\pi^{\mathsf{ref}}(a_{1})=1/C$, where $C-1=e^{\eta}$. A direct computation yields that
$\pi^{*}(a_{3})=\frac{e^{\eta}}{e^{\eta}+C-1},\quad\Rightarrow\quad C^{\pi^{*}}=C\frac{e^{\eta}}{e^{\eta}+C-1}=\frac{C}{2}.$
On the other hand, we know that for $i=1,2$, we have $D^{2}(a_{i})\leq\pi^{\mathsf{ref}}(a_{i})^{-1}\leq 4$. As for $a_{3}$, since we have $\left\langle\bm{\theta},\phi(a_{3})\right\rangle^{2}=\left\langle\bm{\theta},\phi(a_{1})+\phi(a_{2})\right\rangle^{2}\leq 2\left\langle\bm{\theta},\phi(a_{1})\right\rangle^{2}+2\left\langle\bm{\theta},\phi(a_{2})\right\rangle$, which gives that $D^{2}(a_{3})\leq 2D^{2}(a_{1})+2D^{2}(a_{2})\leq 16$. Therefore, taking expectation over $\pi^{*}$, we know that $D^{2}_{\pi^{*}}\leq 12$ which is a constant. ∎
Appendix C Experimental Details
C.1 Linear Bandits
The linear bandit instance used for Figure 2(b) has $d=20$ and $|\mathcal{A}|=100$. For each arm $a\in\mathcal{A}$, we randomly generate its feature vector $\bm{\phi}(a)\in\mathbb{R}^{d}$ such that $\|\bm{\phi}(a)\|=1$. We then randomly sample the model parameter $\bm{\theta}^{*}\in\mathbb{R}^{d}$ such that $\|\bm{\theta}^{*}\|=1$ and the expected reward is obtained via $r(a)=\left\langle\bm{\theta}^{*},\bm{\phi}(a)\right\rangle$.
C.2 Real-Wold Experiments
MNIST consists of 60000 figures, each of which is of $28\times 28$ pixels and consists of a handwritten digit in $\{0,\cdots,9\}$. Here, we consider each image as a context and $\mathcal{A}=\{0,\cdots,9\}$ for each context. To obtain the feature $\bm{\phi}(s,a)$, we first use the hidden representation of a classifier to embed each image as a vector in $\mathbb{R}^{10}$. We then follow the approach in Zhou et al. (2020) to obtain the feature of each context-action pair by having $\bm{\phi}(s,a)=\mathbf{x}\otimes\mathbf{e}_{a+1}\in\mathbb{R}^{100}$, where $\mathbf{x}$ is the output of image encoder and $\otimes$ stands for tensor product.
Appendix D Missing Proofs from Section 2
D.1 Proof of Lemma 2.9